OpenStudy (anonymous):

A 15 foot ladder is leaning against a house with its bottom 9 feet from the base of the house. What is the slope of the ladder?

OpenStudy (anonymous):

need help

OpenStudy (anonymous):

This is a standard right triangle problem. The formula for solving the parameter of a triangle is a^2 + b^2 = c^2 Also known as the Pythagorean theorem. So, in this problem, the ladder is the hypotenuse, that is to say side c of the triangle. The height of the house is 12 or b in this example. Finally, the distance from the house, is the unknown distance a . Applying this knowledge gives us a^2 + 12^2 = 15^2 Simplified to a^2 + 144 = 225 225 - 144 = 81 So now you have a^2 = 81. Taking the square root of each side of the equation, we're left with: a = 9 So, the ladder is nine feet away from the house at the bottom of the ladder Read more: http://wiki.answers.com/Q/A_15feet_ladder_leaning_against_a_bulding_touches_the_wall_12_feet_above_the_ground._How_far_from_the_building_is_the_bottom_of_the_ladder#ixzz2TPN1VHs3

OpenStudy (anonymous):

so that's the slope